OpenStudy (anonymous):

(2k, 3), (1, k); m = 2 Find K

OpenStudy (anonymous):

\(\large 2=\frac{3-k}{2k-1} \) \(\large 2(2k-1)=3-k \) \(\large 4k-2=3-k \) \(\large 5k=5 \) \(\large k=1 \)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw